Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ (a) mean,​ (b) median,​ (c) mode, and​ (d) midrange for the given sample data in millions of dollars. Given that these are the top 10​ salaries, do we know anything about the salaries of TV personalities in​ general? Are such top 10 lists valuable for gaining insight into the larger​ population?

37.7, 35.8, 35.6, 26.9,14.1, 12.8, 12.1, 9.9, 10.3, 8.4

what is the mean ?
Answer #1

here putting above data in ascending order:

S.No X
1 8.40
2 9.90
3 10.30
4 12.10
5 12.80
6 14.10
7 26.90
8 35.60
9 35.80
10 37.70
Σx 203.6
Xbar=Σx/n 20.3600

a)mean =20.36

b) median =middle value =(10+1)/2 th value =5.5 th value =(12.8+14.1)/2=13.45

c) mode=not exist ; as each value appears once so every value is mode or none is,

d)

midrange =(max+min)/2 =(8.4+37.7)/2=23.05

as data is not random and only consider highest 10 values ; therefore it does not represent the whole population.

It is not valuable for gaining insight into the larger​ population